# Bridge Sampling Diagnostics

### v1

@@ 4 Pareto-$\hat{k}$ Diagnostic
$M\;=\;\Bigl{\lfloor}\min\{\,0.2\,S,\;3\sqrt{S}\}\Bigr{\rfloor},\qquad u=Z_{(M)},\qquad E_{i}=Z_{(i)}-u,\;i=1,\dots,M.$
The excesses $\{E_{i}\}$ are modeled with a generalized Pareto distribution
$p(x\mid u,\sigma,k)=\begin{cases}\dfrac{1}{\sigma}\left(1+k\,\dfrac{x}{\sigma}\right)^{-\left(\tfrac{1}{k}+1\right)},&k\neq 0,\;x\geq 0,\\[10.0pt] \dfrac{1}{\sigma}\exp\bigl{(}-x/\sigma\bigr{)},&k=0,\;x\geq 0,\end{cases}$
and the shape parameter $k$ is estimated using the fast approximate Bayesian computation proposed by Zhang and Stephens, (2009), yielding the scalar diagnostic $\hat{k}$. A value
$\hat{k}<0.5\quad\bigl{(}\text{or }0.7\text{ with smoothing}\bigr{)}$
indicates that the empirical mean $\bar{Z}$ behaves well in finite samples ($\hat{k}<0$ corresponding to a bounded distribution), whereas $\hat{k}$ above those thresholds signals behavior typical for heavy-tailed distributions. As the distributions are known to be bounded, increasing the sample sizes $S_{1}$ and $S_{2}$ will eventually lead to $\hat{k}<0$, although that may require impractically large sample sizes.
[…]
$\hat{k}_{N}\;=\;\hat{k}\bigl{(}\{N_{i}\}\bigr{)},\qquad\hat{k}_{D}\;=\;\hat{k}\bigl{(}\{D_{j}\}\bigr{)}.$
This procedure implicitly assumes tail-independence between $N_{i}$ and $D_{j}$, which is true conditionally on the final bridge function after the final iteration as the draws from the posterior and the proposal are independent.
Empirical results in Section 6 show that the component-wise $\hat{k}$ values are informative for bridge sampling and are thus a useful, but not exhaustive safeguard, as they reliably flag many heavy-tail pathologies in bridge sampling. Still, practical usefulness is limited as the tail shape can be far from Pareto shaped.
@@ 5 Block Reshuffling Diagnostic
The MCSE estimate discussed in Section 3 does not take into account potential extra variation due to two algorithm parts: 1) the original posterior draws are split in half with the second half used to estimate the parameters of the proposal distribution, and 2) the potential non-convergence of the iterative updates. The latter can happen if the proposal distribution happens to be bad and the empirical mean $\bar{N}$ or $\bar{D}$ is unstable.
To take into account the effect of these algorithm parts we propose to use a bootstrap style approach. Instead of regular bootstrap, the original posterior draws are divided in blocks and the blocks are shuffled before dividing the posterior draws in half and continuing with the algorithm as usual. The rationale behind using block reshuffling, as opposed to simply random shuffling of the draws, lies in the preservation of this autocorrelation structure, which influences the magnitude of Monte Carlo error. Re-shuffling and re-running bridge sampling $R>1$ times, we get a set of (log) marginal likelihood estimates $\{\hat{p}_{r}(\mathbf{y}\mid\mathcal{M})\}_{r=1}^{R}$. Standard deviation of these can be used as Monte Carlo standard error estimate.
Compared to actually rerunning MCMC, block re-shuffling is expected to be slightly optimistic as the reshuffled sets are overlapping and thus have dependency. Experimental results confirm usefulness of the block reshuffling and the slight optimistic bias.
The Pareto-$\hat{k}$ diagnostic can also be used to diagnose the tail shape of the distribution $\hat{p}_{r}(\mathbf{y}\mid\mathcal{M})$}. Based on the experiments in Section 6, Pareto-$\hat{k}$ for the block reshuffling estimates is more reliable than directly for the bridge sampling numerator and denominator, but has the additional cost of running bridge sampling several times.
[…]
@@ 6 Experimental Results
This section presents the empirical evaluation of the bridge sampling method, along with the proposed diagnostics. We begin by introducing a simple toy example, which serves to illustrate the conditions under which the bridge sampler performs effectively, as well as the circumstances in which it has too high variance to be useful. Following this, we present results using many real models and datasets sourced from the posterior database posteriordb, introduced by Magnusson et al., (2025).
@@ 6.1 Sparse Linear Regression with Increasing Number of Covariates
The simulated data for this analysis were generated as follows. Given $n=100$ observations, the design matrix $\mathbf{X}$ was created with $k$ covariates, where $k$ varies from 10 to 103. Each element of $\mathbf{X}$ was drawn from a standard normal distribution and scaled by $\frac{1}{\sqrt{n\times k}}$:
$\mathbf{X}=\frac{1}{\sqrt{n\times k}}\mathbf{Z},\quad\text{where}\quad\mathbf{Z}\sim\mathcal{N}(0,1)^{n\times k}.$
@@
$\mathbf{y}=\mathbf{X}\boldsymbol{\beta}+\frac{1}{\sqrt{n}}\boldsymbol{\epsilon},\quad\text{where }\boldsymbol{\epsilon}\sim\mathcal{\mathcal{missing}}{N}(0,2).$
@@
For the experiment, it doesn’t matter that the prior does not match how the data is generated. The important aspect is to make the posterior challenging for bridge sampling using a normal proposal distribution.
@@
The posterior sampling was done using the rstanarm function stan_glm (Goodrich et al., , 2024). Bridge sampling was done using bridge_sampler function in the bridgesampling package (Gronau et al., , 2020). We sampled 4000 posterior draws from each posterior, which is the default in rstanarm.
@@
Figure 1 shows the standard deviations of log marginal likelihood and marginal likelihood estimates, and the MCSE estimate for the first. As the number of covariates increases, the posterior gets more high-dimensional and non-normal, increasing the variability of the bridge sampling estimates of the (log) marginal likelihood. The MCSE estimate starts to fail around 50 covariates. Standard deviation estimation via block reshuffling slightly underestimates the standard deviation after about 60 covariates.
@@
Figure 1: Top plot shows the standard deviation of the log marginal likelihood estimate (in log scale) from repeated running of $MC$ and bridge sampling (blue), the standard deviation of log marginal likelihood estimate from repeated running of bridge sampling with block reshuffling (green), and the MCSE estimate from a single bridge sampling estimate (red). The MCSE estimate starts to fail after about 45 covariates. The bottom plot shows the standard deviation (normalized by the median marginal likelihood given the number of covariates) of the marginal likelihood estimate (in log scale) from repeated running of $MC$ and bridge sampling (blue) and the standard deviation of the log marginal likelihood estimate from repeated running of bridge sampling with block reshuffling (green). Standard deviation estimation via block reshuffling slightly underestimates the standard deviation after about 60 covariates. The correspondence between the standard deviation for the log marginal likelihood and the standard deviation for the marginal likelihood is very close to the correspondence between the standard deviations of normal and log normal distributions.
@@
To give some idea of the scale of the potential error, we can consider using the (log) marginal likelihood for model comparison. Based on the experiments, the uncertainty in the log marginal likelihood estimates is quite well approximated by a normal distribution. With bridge sampling, the (log) marginal likelihood estimates for two models are independent, and if the MCSE are similar, then MCSE for difference of log marginal likelihoods is $\sqrt{2}$ times as big. If the true difference would be 0, then a log marginal likelihood MCSE of about 0.5 would lead to 6% of the repeated estimates to be smaller than 1/3 and 6% to be bigger than 3, corresponding to moderate evidence for either model depending on chance. Thus, a desirable MCSE for the log marginal likelihood would be less than 0.2 if the difference is expected to be small (naturally, this depends on the specific use case).
@@
If using Stan for the posterior inference, we recommend to start with the default number of posterior draws (4000), and then adjust based on the observed MCSE. For simple cases, MCSE with the default number of draws may already be sufficient [^1] At the time of writing, the bridgesampling package documentation states “The computation of marginal likelihoods based on bridge sampling requires a lot more posterior draws than usual. A good conservative rule of thump is perhaps 10-fold more draws.”. If the estimated MCSE is too high, more posterior draws can be obtained. As the variance of the estimator is finite, we may assume in many cases that MCSE halves by quadrupling the number of posterior draws and use that rule to estimate how many posterior draws are required for the desired accuracy. If the bound for either numerator or denominator in the bridge sampling equation is far, then even more draws may be required for the desired accuracy (Vehtari et al., , 2024).
@@
In this example the biggest standard deviation for the log marginal likelihood is 2.5, which in comparing models with no true difference would lead to 10% of the repeated estimates to be smaller than 1/100 and 10% to be bigger than 100, corresponding to extreme evidence for either model depending on luck. To get a MCSE of 2.5 reduced to around 0.2 would require at least 156 times more posterior draws, that is, more than 600 000 posterior draws (and might require more if the pre-asymptotic convergence rate is lower due a long tail and very large bound). Alternate proposal distributions may also improve the performance.
@@
We next look at the use of the Pareto-$\hat{k}$ as a further diagnostic. Figure 2 shows the average Pareto-$\hat{k}$ diagnostic using the default number of the tail values (135 given 2000 draws, continuous lines). Pareto-$\hat{k}$s for both numerator (yellow) and denominator (blue) are below 0.5 when the number of covariates is less than 21, and thus, in those cases, we could trust MCSE estimates.
@@
Figure 2: Pareto-$\hat{k}$ diagnostic in the linear regression example with the increasing number of covariates for the distribution of terms in the numerator (yellow) and denominator (blue). The plot shows the average over 100 independent MCMC runs of estimated $\hat{k}$ using the default number of tail draws (135) and only 10 tail draws with continuous and dashed lines, respectively.
@@
However, Figure 1 did show that the MCSE estimate was accurate up to about 45 covariates. In this case, the tails of the numerator and denominator distributions are far from Pareto shape, and Pareto-$\hat{k}$ with the default number of tail values gives too high values. Figure 2 also shows the average Pareto-$\hat{k}$ diagnostic using only 10 tail values (dashed lines). Now Pareto-$\hat{k}$s for both numerator and denominator are below 0.5 when the number of covariates is less than 42, which better matches the range when the MCSE estimate starts to fail. This shows that even when the distributions are bounded, with a finite sample the tail draws can resemble draws from a thick tail leading to an unreliable variance-based MCSE estimate. In this case, the average Pareto-$\hat{k}$ diagnostic using only 10 tail values matches the average behavior, but for a single case the Pareto-$\hat{k}$ using only 10 tail values has too big variability to be a practical diagnostic. We leave it for future research to develop an alternative tail shape diagnostic that would work well for such bounded distributions that arise in bridge sampling.
@@
To gain further insights into the performance of bridge sampling under different conditions, we proceed to an examination of 41 models and posteriors from the posteriordb database (Magnusson et al., , 2025) plus a multi-component Gaussian process model of birthday data (Zhang et al., , 2022). The posterior distributions selected for this study are the same as those analyzed by Zhang et al., (2022) plus several additional high dimensional posteriors that have been included in posteriordb more recently. They include a representative set of real life posteriors with varying dimensionality, non-normality, funnelness, and multi-modality. The list of posteriors, the number of posterior dimensions (in unconstrained space), and short description of the model is provided in Appendix A. For many models the Stan code used is dropping the constant terms, so the estimated marginal likelihoods are also missing that constant, but that does not affect the estimation variability. We sampled 4000 posterior draws from each posterior using CmdStanR with the default options except using the Pathfinder algorithm (Zhang et al., , 2022) to initialize MCMC for improved convergence speed and reliability. We did check that the MCMC sampling performed reasonably for all posteriors, but in some cases there are divergences which indicate the possibility of biased estimates, which can also increase the bias and variability in the (log) marginal likelihood estimates. We did not compute true marginal likelihood for any of the posteriors, and it would be impossible for most of the posteriors, but compare the performance to the average over 100 independent repeated runs.
@@
Figure 3 shows MCSE and standard deviation of log marginal likelihood estimates from block reshuffling vs standard deviation of log marginal likelihood estimates from repeated MCMC runs. MCSE estimates are quite accurate for most of the posteriors while block reshuffling standard deviation tends to be slightly smaller than the MCMC standard deviation when the true standard deviation is also small. For 36/42 posteriors MCSE and MCMC standard deviation are below 0.2. For the Birthdays posterior we can estimate that 13 times more posterior draws, that is, at least 52 000 posterior draws would be required to get a MCSE of 0.2 or smaller. For the Presidential votes hierarchical Gaussian process posterior we can estimate that 500 times more posterior draws, that is, at least 2 million posterior draws would be required to get MCSE of 0.2 or smaller, which starts to be impractically big sample size.
@@
Figure 3: posteriordb posteriors + Birthdays: The left plot shows estimated MCSE vs standard deviation of log marginal likelihood estimates from repeated MCMC runs. The right plot shows standard deviation of log marginal likelihood estimates from block reshuffling vs standard deviation of log marginal likelihood estimates from 100 repeated MCMC runs. MCSE estimate has small bias except for a few posteriors with the highest variability. Block reshuffling underestimates the small variability estimates, but has small bias in case of high variability.
@@
Figure 4: posteriordb posteriors + Birthdays: The number of posterior dimensions vs standard deviation of log marginal likelihood estimates from repeated MCMC runs. The estimate variability tends to increase with the number of dimensions, which is natural due to the curse of dimensionality. In addition there is variation depending on how non-normal the posterior distribution is.

### v2

@@ 4 Pareto-$\widehat{k}$ Diagnostic and Pre-asymptotic Stability
$0\leq N_{i}\leq\frac{1}{s_{1}},\qquad 0\leq D_{j}\leq\frac{1}{s_{2}}\quad\text{a.s.}$ (11)
Consequently, all moments of $N_{i}$ and $D_{j}$ are theoretically finite, and the central limit theorem (CLT) holds asymptotically. However, asymptotic finite variance is insufficient for reliability in practice, as the convergence of the MCSE estimator depends on the sample variance converging to the population variance. We identify a specific regime where this convergence is critically slow, rendering standard variance estimators unreliable.
Define the pre-asymptotic regime $\mathcal{R}_{\text{pre}}$ as the set of posterior-proposal pairs $(\pi,g)$ and sample sizes $S_{1}$ and $S_{2}$, where the effective support of the bridge function is unexplored. Let $r(\mathrm{\boldsymbol{\theta}})=\widetilde{\pi}(\mathrm{\boldsymbol{\theta}})/g(\mathrm{\boldsymbol{\theta}})$ be the unnormalized importance ratio. We say that a bridge sampling estimation is in $\mathcal{R}_{\text{pre}}$ if the proposal $g(\mathrm{\boldsymbol{\theta}})$ fails to uniformly cover the typical set of $\pi(\mathrm{\boldsymbol{\theta}})$, such that the distribution of the importance weights exhibits heavy-tailed behavior within the observable range of $S_{1}$ and $S_{2}$. This occurs under two primary conditions:
  1. (i)
    Tail Mismatch: The target $\pi(\mathrm{\boldsymbol{\theta}})$ has heavier tails than the proposal $g(\mathrm{\boldsymbol{\theta}})$ (e.g., a Student-$t$ posterior approximated by a Gaussian proposal); here, $r(\mathrm{\boldsymbol{\theta}})\to\infty$ in the tails, and while $N_{i}$ is bounded by $1/s_{1}$, the probability of sampling near this bound is negligible for realistic $S_{2}$.
  2. (ii)
    High-Dimensional Mismatch: Even with matching tail types, if $D$ is sufficiently large, the variance of the log-weights satisfies $\operatorname{Var}_{g}[\log r(\mathrm{\boldsymbol{\theta}})]\gg 1$. The distribution of $r(\mathrm{\boldsymbol{\theta}})$ becomes approximately Log-Normal with high variance, indistinguishable from a heavy-tailed distribution in finite samples.
In $\mathcal{R}_{\text{pre}}$, the values of $N_{i}$ and $D_{j}$ follow a power-law distribution $\mathbb{P}[X>x]\propto x^{-1/k}$ before truncating at the theoretical bounds. Standard variance estimators fail here as they are sensitive to the unobserved frequency of events near the bounds (Vehtari et al.,, 2024).
We propose the Pareto-$\widehat{k}$ diagnostic as a necessary condition for sample size sufficiency. We fit the generalized Pareto distribution to the upper tails of $\{N_{i}\}$ and $\{D_{j}\}$:
$\widehat{k}_{N}=\widehat{k}(\{N_{i}\}),\qquad\widehat{k}_{D}=\widehat{k}(\{D_{j}\}).$
Following the convergence rate analysis of importance sampling ratios in Vehtari et al., (2024), we establish the following validity criteria based on the estimated tail index $\widehat{k}$:
  1. (i)
    Finite variance regime ($\widehat{k}\leq 0.5$): When the tail index $k\leq 0.5$, the underlying distribution has a finite variance. In this regime, the standard central limit theorem applies, and the Monte Carlo error decreases at the standard rate of $\mathcal{O}(S^{-1/2})$. Consequently, the standard MCSE (Section 3) is reliable and correctly characterizes the uncertainty of the estimator.
  2. (ii)
    Smoothing-indicated regime ($0.5<\widehat{k}\leq 0.7$): In this range, the weight distribution exhibits pre-asymptotic behavior consistent with infinite variance, meaning the standard central limit theorem does not hold. In standard importance sampling, this instability is effectively mitigated using Pareto Smoothed Importance Sampling (PSIS, Vehtari et al.,, 2024), however, in bridge sampling, the terms $N_{i}$ and $D_{j}$ are strictly bounded (Equation 11), and empirically we observe that the tail index estimator $\widehat{k}$ is often positively biased in this setting, interpreting the distribution’s slow decay toward its bound as a heavy, unbounded tail. Thus, applying Pareto smoothing based on an overestimated $\widehat{k}$ risks imputing values that violate these bounds or assigning excessive probability mass to the tails, thereby potentially biasing the marginal likelihood estimate. Consequently, we do not apply smoothing and retain the raw estimator, which, while usable, exhibits convergence rates slower than $\mathcal{O}(S^{-1/2})$, rendering the MCSE a likely underestimate of the true uncertainty.
  3. (iii)
    Failure warning ($\widehat{k}>0.7$): If $\widehat{k}>0.7$, the estimator operates deeply in the pre-asymptotic regime. The effective sample size is insufficient to explore the typical set, and the estimate is dominated by rare events. In this context, the variance-based MCSE is unreliable, and the result should not be trusted without increasing the sample size or improving the proposal.
When $\widehat{k}>0.5$, practitioners should exercise caution, and while values up to $0.7$ may be tolerated if one accepts that the reported MCSE is optimistic, a high $\widehat{k}$ suggests increasing the sample size $S$ significantly or utilizing the Block Reshuffling diagnostic (Section 5) to obtain a more robust empirical estimate of the uncertainty.
@@ 5 Block Reshuffling Diagnostic
The analytic MCSE derived in Section 3 relies on the asymptotic normality guaranteed by the central limit theorem; however, as discussed in Section 4, in the pre-asymptotic regime the distribution of the bridge sampling estimator is often heavy-tailed or skewed, and in such cases, the symmetric confidence intervals implied by the standard error are often misleading. To diagnose these finite-sample pathologies, we propose a block reshuffling diagnostic, which adapts the blocking strategy of the moving block bootstrap (Kunsch,, 1989) to the specific workflow of bridge sampling, using permutations to assess stability.
Standard resampling methods assume independent and identically distributed data; however, as MCMC draws exhibit autocorrelation, simple reshuffling destroys the dependence structure necessary for valid inference; the block approach addresses this by resampling contiguous blocks of the Markov chain, thereby preserving the local dependence structure (autocorrelation) within blocks while treating the blocks themselves as exchangeable (Liu and Singh,, 1992).
Our procedure operates as follows: let $\mathcal{S}=(\mathrm{\boldsymbol{\theta}}^{*}_{1},\dots,\mathrm{\boldsymbol{\theta}}^{*}_{S})$ denote the available posterior draws, we partition $\mathcal{S}$ into $B$ contiguous blocks; to generate a single replicate $\widehat{Z}_{r}$, we resample these blocks without replacement (i.e., permute the blocks) to form a new sequence of length $S$, and we then run the full bridge sampling procedure on this reshuffled sequence: the draws are split, the proposal $g(\mathrm{\boldsymbol{\theta}})$ is fitted to the first half, and the iterative update (Equation 5) is run to convergence. This process is repeated $R$ times to generate an empirical distribution of estimates $\{\widehat{Z}_{r}\}_{r=1}^{R}$.
By re-estimating the proposal density for each replicate, this procedure approximates the sampling distribution of the estimator conditional on the empirical distribution of the chain. Unlike the analytic MCSE, which is a moment-based estimator fixed to a single proposal fit, the reshuffling ensemble captures the sensitivity of the estimator to the specific partitioning of the posterior draws and the resulting proposal approximation.
We define the block reshuffling error as the standard deviation of the log-transformed replicates:
$\widehat{\mathrm{MCSE}}_{\mathrm{BR}}(\log\widehat{Z})\coloneqq\sqrt{\frac{1}{R-1}\sum_{r=1}^{R}\left(\log\widehat{Z}_{r}-\overline{\log\widehat{Z}}\right)^{2}}.$
More importantly, the bootstrap distribution allows us to assess properties beyond variance, specifically skewness and tail behavior. As noted by Hall, (1992), resampling methods are particularly effective at characterizing the shape of a sampling distribution, providing a more accurate picture of uncertainty than CLT-based approximations.
Lastly, we explicitly apply the Pareto-$\widehat{k}$ diagnostic to the replicates $\{\widehat{Z}_{r}\}_{r=1}^{R}$ as a stability check. Empirically, this aggregate diagnostic proves more stable than monitoring the raw numerator and denominator contributions (Section 6). A high Pareto-$\widehat{k}$ value in this setting serves as a supplementary warning signal, implying that the estimator’s instability is structural; in such cases, even repeated sampling from the posterior cannot guarantee reliability without further computational investment.
We caution, however, that while block reshuffling detects non-normality, it likely underestimates the total variance. Because the procedure permutes a fixed, finite realization of the Markov chain rather than sampling from the true posterior, it cannot fully capture the variation that would arise from independent MCMC runs (Gelman and Rubin,, 1992). Consequently, $\widehat{\mathrm{MCSE}}_{\mathrm{BR}}$ should be viewed as a diagnostic lower bound: if the reshuffling error is large or the distribution is skewed, the estimate is certainly unstable; however, a small reshuffling error does not guarantee accuracy if the original chain failed to explore the typical set.
@@ 6 Experimental Results
This section presents the empirical evaluation of the bridge sampling method, along with the proposed diagnostics. Our analysis focuses primarily on the precision (variability) of the estimates rather than systematic bias, as we utilize the standard bridge sampling estimator without bias-inducing modifications (even in the pre-asymptotic regime, see Section 4), which implies that the estimator remains consistent and asymptotically unbiased. In the finite-sample regime, errors are driven by variance and skewness rather than systematic shift; thus, assessing the reliability of the MCSE is equivalent to assessing the root mean square error (RMSE).
As analytical ground truth is unavailable for the complex posteriors in posteriordb, we establish a computational “gold standard” by aggregating 100 independent MCMC and bridge sampling runs (totaling 400,000 posterior draws with low autocorrelation per model), which will serve as the proxy for the true marginal likelihood against which individual MCSE predictions are validated.
We begin by introducing a simple toy example, which serves to illustrate the conditions under which the bridge sampler performs effectively, as well as the circumstances in which it has too high variance to be useful. Following this, we present results using many real models and datasets sourced from the posterior database posteriordb, introduced by Magnusson et al., (2025).
@@ 6.1 Sparse Linear Regression with Increasing Number of Covariates
The simulated data for this analysis were generated as follows. Given $n=100$ observations, the design matrix $\mathbf{X}$ was created with $k$ covariates, where $k\in\{10,11,12,\dots,103\}$. Each element of $\mathbf{X}$ was drawn from a standard normal distribution and scaled by $\frac{1}{\sqrt{n\times k}}$:
$\mathbf{X}=\frac{1}{\sqrt{n\times k}}\boldsymbol{\xi},\quad\text{where}\quad\boldsymbol{\xi}\sim\mathcal{N}(0,1)^{n\times k}.$
@@
$\mathbf{y}=\mathbf{X}\boldsymbol{\beta}+\frac{1}{\sqrt{n}}\boldsymbol{\epsilon},\quad\text{where }\boldsymbol{\epsilon}\sim\mathcal{N}(0,2).$
@@
This mismatch between the data generation (Normal) and the estimation model (RHS) is intentional; it creates a challenging, funnel-like posterior geometry for the bridge sampler to integrate.
@@
The posterior sampling was done using the rstanarm function stan_glm (Goodrich et al.,, 2024). Bridge sampling was done using bridge_sampler function in the bridgesampling package (Gronau et al., 2020a, ). We sampled 4000 posterior draws from each posterior, which is the default in rstanarm.
@@
Figure 2 shows the standard deviations of log marginal likelihood and marginal likelihood estimates, and the MCSE estimate for the first. As the number of covariates increases, the posterior gets more high-dimensional and non-normal, increasing the variability of the bridge sampling estimates of the (log) marginal likelihood, with the analytic MCSE estimate starting to lose precision around 50 covariates. Consistent with our discussion in Section 5, the block reshuffling diagnostic slightly underestimates the total standard deviation in the highest dimensions due to the reuse of draws, yet correctly identifies the regime of instability. Furthermore, the correspondence between the standard deviation for the log marginal likelihood and the standard deviation for the marginal likelihood is very close to the correspondence between the standard deviations of normal and log normal distributions.
@@
Figure 2: Top plot shows the standard deviation of the log marginal likelihood estimate (in log scale) from repeated running of MCMC and bridge sampling (solid line), the standard deviation of log marginal likelihood estimate from repeated running of bridge sampling with block reshuffling (long dashed line), and the MCSE estimate from a single bridge sampling estimate (dashed line). The bottom plot shows the standard deviation (normalized by the median marginal likelihood given the number of covariates) of the marginal likelihood estimate (in log scale) from repeated running of MCMC and bridge sampling (solid line) and the standard deviation of the log marginal likelihood estimate from repeated running of bridge sampling with block reshuffling (long dashed line).
@@
To give some idea of the scale of the potential error, we can consider using the (log) marginal likelihood for model comparison. Based on the experiments, the uncertainty in the log marginal likelihood estimates is quite well approximated by a normal distribution. With bridge sampling, the log marginal likelihood estimates for two models are independent, and if the MCSE are similar, then MCSE for difference of log marginal likelihoods is $\sqrt{2}$ times as big. If the true difference were 0, then a log marginal likelihood MCSE of about 0.5 would lead to 6% of the repeated estimates to be smaller than 1/3 and 6% to be bigger than 3, corresponding to moderate evidence for either model solely due to Monte Carlo variability. Thus, a desirable MCSE for the log marginal likelihood would be less than 0.2 if the difference is expected to be small. Naturally, this depends on the specific use case.
@@
If using Stan for the posterior inference, we recommend starting with the default number of posterior draws (4000), and then adjusting based on the observed MCSE. For simple cases, MCSE with the default number of draws may already be sufficient (at the time of writing, the bridgesampling package documentation states “The computation of marginal likelihoods based on bridge sampling requires a lot more posterior draws than usual. A good conservative rule of thumb is perhaps 10-fold more draws”). If the estimated MCSE is too high, more posterior draws can be obtained. As the variance of the estimator is finite, we may assume in many cases that MCSE halves by quadrupling the number of posterior draws (based on CLT) and use that rule to estimate how many posterior draws are required for the desired precision. If the bound for either numerator or denominator in the bridge sampling equation is far, then even more draws may be required for the desired precision (Vehtari et al.,, 2024).
@@
In this example the biggest standard deviation for the log marginal likelihood is 2.5, which in comparing models with no true difference would lead to 10% of the repeated estimates to be smaller than 1/100 and 10% to be bigger than 100, corresponding to extreme evidence for either model depending on the specific realization of the Monte Carlo draws. To get a MCSE of 2.5 reduced to around 0.2 would require at least 156 times more posterior draws, that is, more than 600 000 posterior draws (and might require more if the pre-asymptotic convergence rate is lower due to a long tail and very large bound). Alternate proposal distributions may also improve the performance.
@@
We next look at the use of the Pareto-$\widehat{k}$ as a further diagnostic. Figure 3 shows the average Pareto-$\widehat{k}$ diagnostic using the default number of the tail values (135 given 2000 draws, continuous lines). Pareto-$\widehat{k}$’s for both numerator and denominator are below 0.5 when the number of covariates is less than 21, and thus, in those cases, we could trust MCSE estimates.
@@
Figure 3: Pareto-$\widehat{k}$ diagnostic in the linear regression example with the increasing number of covariates for the distribution of terms in the numerator and denominator (as labeled in the plot). The plot shows the average over 100 independent MCMC runs of estimated $\widehat{k}$ using the default number of tail draws (135) and only 10 tail draws with continuous and long dashed lines, respectively.
@@
However, Figure 2 did show that the MCSE estimate was precise up to about 50 covariates. In this case, the tails of the numerator and denominator distributions are far from Pareto shape, and Pareto-$\widehat{k}$ with the default number of tail values gives too high values. Figure 3 also shows the average Pareto-$\widehat{k}$ diagnostic using only 10 tail values (dashed lines). Now Pareto-$\widehat{k}$s for both numerator and denominator are below 0.5 when the number of covariates is less than 42, which better matches the range when the MCSE estimate starts to fail. This confirms that even though the distributions are theoretically bounded, in the pre-asymptotic regime the finite-sample draws exhibit heavy-tailed behavior indistinguishable from unbounded distributions, rendering the variance-based MCSE unreliable. In this case, the average Pareto-$\widehat{k}$ diagnostic using only 10 tail values matches the average behavior, but for a single case the Pareto-$\widehat{k}$ using only 10 tail values has too big variability to be a practical diagnostic. We leave it for future research to develop an alternative tail shape diagnostic that would work well for such bounded distributions that arise in bridge sampling.
@@
To gain further insights into the performance of bridge sampling under different conditions, we proceed to an examination of 41 models and posteriors from the posteriordb database (Magnusson et al.,, 2025) plus a multi-component Gaussian process model of birthday data (Zhang et al.,, 2022). The posterior distributions selected for this study are the same as those analyzed by Zhang et al., (2022) plus several additional high-dimensional posteriors that have been included in posteriordb more recently. They include a representative set of real-life posteriors with varying dimensionality, non-normality, funnelness, and multi-modality. The list of posteriors, the number of posterior dimensions (in unconstrained space), and a short description of the model are provided in Appendix A. For many models the Stan code used is dropping the constant terms, so the estimated marginal likelihoods are also missing that constant, but that does not affect the estimation variability. We sampled 4000 posterior draws from each posterior using CmdStanR with the default options except using the Pathfinder algorithm (Zhang et al.,, 2022) to initialize MCMC for improved convergence speed and reliability. We did check that the MCMC sampling performed reasonably for all posteriors, but in some cases there are divergences which indicate the possibility of biased estimates, which can also increase the bias and variability in the (log) marginal likelihood estimates. We did not compute true marginal likelihood for any of the posteriors, and it would be impossible for most of the posteriors, but compare the performance to the average over 100 independent repeated runs. To quantify the computational gain, for one of the most complex models in our set (Birthdays GP), these 100 independent MCMC runs took 1571.4 minutes, whereas generating 100 estimates via block reshuffling took only 72.8 minutes (a $\approx 21.5\times$ speedup).
@@
Figures 4 and 5 show MCSE and standard deviation of log marginal likelihood estimates from block reshuffling vs standard deviation of log marginal likelihood estimates from repeated MCMC runs. MCSE estimates are quite precise for most of the posteriors while block reshuffling standard deviation tends to be slightly smaller than the MCMC standard deviation when the true standard deviation is also small. For 36/42 posteriors MCSE and MCMC standard deviation are below 0.2. For the Birthdays posterior we can estimate that 13 times more posterior draws, that is, at least 52 000 posterior draws would be required to get a MCSE of 0.2 or smaller. For the Presidential votes hierarchical Gaussian process posterior we can estimate that 500 times more posterior draws, that is, at least 2 million posterior draws would be required to get MCSE of 0.2 or smaller, which starts to be an impractically large number of draws.
@@
Figure 4: posteriordb posteriors + Birthdays: Estimated MCSE vs standard deviation of log marginal likelihood estimates from repeated MCMC runs. MCSE estimate has small bias except for a few posteriors with the highest variability.
Figure 5: posteriordb posteriors + Birthdays: Standard deviation of log marginal likelihood estimates from block reshuffling vs standard deviation of log marginal likelihood estimates from 100 repeated MCMC runs. Block reshuffling underestimates the small variability estimates, but has small bias in case of high variability.
@@
Figure 6: posteriordb posteriors + Birthdays: The number of posterior dimensions vs standard deviation of log marginal likelihood estimates from repeated MCMC runs. The estimate variability tends to increase with the number of dimensions, which is natural due to the curse of dimensionality. In addition, there is variation depending on how non-normal the posterior distribution is.